OpenStudy (anonymous):

Dimensional Analysis Is my reasoning correct? *problem* Oil with a given viscosity and density flows through a pipe with given diameter and flowrate. We want to make a scale model with water that flows through a pipe with a smaller given diameter. I have to find out what the speed should be in the scalemodel to achieve similar viscous forces. *reasoning* variables: v,mu,rho,D units: kg,m,s v = k*(mu^a * rho^b * D^c) (with k being some dimensionless constant) [m/s] = [kg/m*s]^a * [kg/m^3]^b * [m]^c => a+b=0 => a=1 <=> b=-1 => 1=-a-3b+c <=> c=-2 so for the real model v_r

OpenStudy (michele_laino):

If we consider the flowrate and I call with \(\eta\) the viscosity, with \(\delta \) the water density, with \(D\) the diameter of the pipe, and with \(Q\) the flowrate of water. Then I can write this: \[\Large k\; {\eta ^a}{\delta ^b}{D^c}{Q^d} = v\] (where \(k\) is a dimensionless quantity) so, replacing each quantity, with the corresponding unit of measure, I get: \[\Large {\left( {\frac{{Kg}}{{m \cdot s}}} \right)^a}{\left( {\frac{{Kg}}{{{m^3}}}} \right)^b}{m^c}{\left( {\frac{{{m^3}}}{s}} \right)^d} = \frac{{m}}{s}\] Finally, I can write the subsequent system: \[\Large \left\{ \begin{gathered} a + b = 0 \hfill \\ c + 3d - a - 3b = 1 \hfill \\ a + d = 1 \hfill \\ \end{gathered} \right.\] nevertheless it is an undetermined system Whereas if I don't consider the flowrate \(Q\), then I can write: \[\Large \begin{gathered} k\;{\eta ^a}{\delta ^b}{D^c} = v \hfill \\ \hfill \\ {\left( {\frac{{Kg}}{{m \cdot s}}} \right)^a}{\left( {\frac{{Kg}}{{{m^3}}}} \right)^b}{m^c} = \frac{m}{s} \hfill \\ \hfill \\ \left\{ \begin{gathered} a + b = 0 \hfill \\ c - a - 3b = 1 \hfill \\ a = 1 \hfill \\ \end{gathered} \right. \hfill \\ \end{gathered} \]